# ReXplain: Translating Radiology into Patient-Friendly Video Reports

### v1

@@ 4 Eliciting User Feedback
In clinical practice, limited time and heavy workload often prevents radiologists from directly communicating with patients (Kemp et al., 2017). The study by Recht et al. (2022) showed that hand-crafted video reports by radiologists were helpful in improving patients’ understanding. However, this approach offered limited information with increased workload (each video had 55 ± 30 seconds of duration and was generated in 238 ± 141 seconds). In contrary, ReXplain functions end-to-end without requiring extra time or effort from radiologists. Furthermore, the explanation could be tailored to provide flexibility for customization through appropriate prompting. We observed that 66% of responses were optimistic about walking patients through the videos (Question #5), which indicates that the generated video reports effectively conveyed the information we intended for patients to understand. This result further demonstrated that ReXplain holds promise in improving patients understanding and experience with lighter workload on radiologists.
5 Discussion
Current Limitations and Outlook.
This proof-of-concept study aimed to to figure out whether and how the AI-driven pipeline, ReXplain, could help translate the conventional radiology reports into patient-friendly video reports. Nevertheless, our user study was conducted based on radiologists, of whom the feedback may not precisely reflect the experience of real patients. The current work shows that the video reports overall were considered useful for enhancing patients’ understanding, and this finding motivates us to conduct future works involving patients without medical knowledge to further investigate the usefulness of the AI-generated video reports.
@@ 5 Discussion
We acknowledge another noteworthy limitation of this study: merely the organs containing the abnormal findings were highlighted, which is limited by the current state of medical image segmentation. A significant portion of written feedback from the radiology team suggested that highlighting entire organs may have limited effectiveness in improving understanding, particularly for small findings such as lung nodules. This probably led to predominantly negative feedback regarding the usefulness of organ rendering for patient comprehension (question #11) and that only 32% of responses (question #6) indicated acceptance of allowing patients to view the videos without supervision. We believe this challenge can be addressed with advancements in text-prompted lesion segmentation, such as BiomedParse (Zhao et al., 2024). We are also interested in exploring whether and how more accurate identification and localization of findings might alter users’ opinions.
@@ Conclusion
In this paper, we present ReXplain, an innovative integration of existing state-of-the-art modules to create a functional, end-to-end system that produces novel, patient-friendly radiology video reports. We showcase an integrative AI pipeline with a practical interface design to solve a complex healthcare communication challenge, bridging a crucial gap in patient understanding of their diagnostic results. The user study by five board-certified radiologists yielded encouraging responses regarding the usefulness of the AI generated video reports. By developing and evaluating ReXplain, this paper lays the groundwork for AI-assisted patient-centered radiology.
Appendix A Prompting GPT-4 for report explanation
A.1 Matching report phrases to organs.
@@ Appendix A Prompting GPT-4 for report explanation
A.2 Report phrase explanation.
@@
Output: $\{$‘‘abnormality\_explanation": ‘‘Your left subclavian vein is found to be collapsed. This vein is an important blood vessel that carries blood from your arm back to your heart. When we say its collapsed, it means that the vein is not as open as it should be, which might affect how blood flows through it.",
‘‘input\_scan\_appearance": ‘‘If we look at your CT image, the left subclavian vein will appear much narrower or even almost closed in some areas. It might look like a thin line instead of a round, hollow tube.",
‘‘normal\_scan\_appearance": ‘‘In a normal CT image, this vein would appear open and round, kind of like a small, hollow tube. It would look uniform and not pinched or narrow."\}
A.3 List of Organs.
@@
right lung lower lobe’, right lung middle lobe’, right lung upper lobe’, lung lower lobe’,
lung upper lobe’, trachea and bronchie’, trachea’, bronchie’, mediastinum’,
brachiocephalic trunk’, brachiocephalic vein’, left brachiocephalic vein’,
@@
spinal cord’, spinal canal’, vertebrae’, cervical vertebrae’, cervical vertebrae 1 (c1)’,
cervical vertebrae 2 (c2)’, cervical vertebrae 3 (c3)’, cervical vertebrae 4 (c4)’,
cervical vertebrae 5 (c5)’, cervical vertebrae 6 (c6)’, cervical vertebrae 7 (c7)’,
thoracic vertebrae’, thoracic vertebrae 1 (t1)’, thoracic vertebrae 2 (t2)’,
thoracic vertebrae 3 (t3)’, thoracic vertebrae 4 (t4)’, thoracic vertebrae 5 (t5)’,
thoracic vertebrae 6 (t6)’, thoracic vertebrae 7 (t7)’, thoracic vertebrae 8 (t8)’,
thoracic vertebrae 9 (t9)’, thoracic vertebrae 10 (t10)’, thoracic vertebrae 11 (t11)’,
thoracic vertebrae 12 (t12)’, lumbar vertebrae’, lumbar vertebrae 1 (l1)’,
lumbar vertebrae 2 (l2)’, lumbar vertebrae 3 (l3)’, lumbar vertebrae 4 (l4)’,
lumbar vertebrae 5 (l5)’, lumbar vertebrae 6 (l6)’, sacral vertebrae 1 (s1)’, clavicle’,
left clavicle’, right clavicle’, scapula’, left scapula’, right scapula’, humerus’,
left humerus’, right humerus’, femur’, left femur’, right femur’, head of femur’,
left head of femur’, right head of femur’, rib’, left rib’, right rib’, rib 1’,
left rib 1’, right rib 1’, rib 2’, left rib 2’, right rib 2’, rib 3’, left rib 3’,
right rib 3’, rib 4’, left rib 4’, right rib 4’, rib 5’, left rib 5’, right rib 5’,
rib 6’, left rib 6’, right rib 6’, rib 7’, left rib 7’, right rib 7’, rib 8’,
left rib 8’, right rib 8’, rib 9’, left rib 9’, right rib 9’, rib 10’, left rib 10’,
right rib 10’, rib 11’, left rib 11’, right rib 11’, rib 12’, left rib 12’,
right rib 12’, rib cartilage’, costal cartilage’, sternum’, manubrium of sternum’,
eustachian tube bone’, left eustachian tube bone’, right eustachian tube bone’, thyroid’,
thyroid gland’, left thyroid’, right thyroid’, breast’, left breast’, right breast’,
abdomen’, abdominal tissue’, adrenal gland’, left adrenal gland’, right adrenal gland’,
colon’, duodenum’, gallbladder’, intestine’, small bowel’, kidney’, left kidney’,
right kidney’, liver’, left lobe of liver’, left lateral inferior segment of liver’,
left lateral superior segment of liver’, left medial segment of liver’, right lobe of liver’,

### v2

@@ Introduction
In clinical practice, limited time and heavy workload often prevents radiologists from directly communicating with patients (Kemp et al. 2017). The study by Recht et al. (2022) showed that hand-crafted video reports by radiologists were helpful in improving patients’ understanding. However, this approach offered limited information with increased workload (each video had 55 ± 30 seconds of duration and was generated in 238 ± 141 seconds). In contrary, ReXplain functions end-to-end without requiring extra time or effort from radiologists, offering enhanced scalability in practice. Furthermore, the explanation could be tailored to provide flexibility for customization through appropriate prompting. We observed that 66% of responses were optimistic about walking patients through the videos (Question #5), which indicates that the generated video reports effectively conveyed the information we intended for patients to understand. This result further demonstrated that ReXplain holds promise in improving patients understanding and experience with lighter workload on radiologists.
Round 2
User Study Design.
Based on feedback from Round 1, we conducted a second round of user feedback to gain deeper insights into the proposed pipeline. We used a structured survey with open-ended questions that explored ReXplain’s limitations, potential improvements, and clinical applications, focusing on exploring features that might discourage patient sharing, addressing safety considerations, and discussing the usefulness of videos without pathology highlighting. Four radiologists participated in this round, including two with 6-10 years of practice and two with 11-15 years of experience. Three of these radiologists had previously completed the Round 1 survey.
Current Limitations and Potential Improvements.
@@
The current organ-level segmentation approach was identified as a key limitation, with radiologists emphasizing the need for more precise pathology highlighting and visualization. A significant portion of written feedback from the radiology team suggested that highlighting entire organs may have limited effectiveness in improving understanding, particularly for small findings such as lung nodules. This probably led to predominantly negative feedback in Round 1 regarding the usefulness of organ rendering for patient comprehension (question #11) and that only 32% of responses (question #6) indicated acceptance of allowing patients to view the videos without supervision. We deem that the limitation is majorly caused by the current state of medical image segmentation models, as few support free-text-based grounding of lesions. While we believe that this challenge can be addressed with advancements in text-prompted lesion segmentation, such as BiomedParse (Zhao et al. 2024). Another limitation pointed is that the continuous video scrolling made it challenging for viewers to focus on specific findings, suggesting a need for strategic pausing at key points. Also, this can be addressed by improving radiology image segmentation to the lesion level, which would enable key slice display focusing on the findings. The feedback also highlighted the importance of avoiding misleading severity assessments and integrating trusted radiology information sources. Improvement was also suggested to implement variable explanation complexity levels and including clear directives for patients to consult healthcare providers.
Potential Clinical Usage.
Building on earlier discussions, ReXplain is envisioned as a pre-consultation educational tool for radiologists. It is suggested to provide ReXplain-generated video reports to patients “probably immediately before the physician sees them.”, followed by “in-person review and question-answer session”, to secure its usage. Moreover, one of the collected suggestions is “these available with every imaging study and look for improvements in patient engagement and follow up, and decreased burden on the referring provider (who ordered the study and typically would have to explain/discuss with the patient)”. While it was emphasized that precise lesion identification would significantly enhance clinical value, three out of four respondents maintained that the videos retain utility even without specific pathology highlighting. These insights reinforce our initial findings while providing concrete directions for technical improvements and safety considerations in clinical deployment.
Outlook.
This proof-of-concept study aimed to to figure out whether and how the AI-driven pipeline, ReXplain, could help translate the conventional radiology reports into patient-friendly video reports. Nevertheless, our user study was conducted by radiologists, of whom the feedback may not precisely reflect the experience of real patients. The current work shows that the video reports overall were considered useful for enhancing patients’ understanding, and this finding motivates us to conduct future works involving patients without medical knowledge to further investigate the usefulness of the AI-generated video reports.
@@ Conclusion
In this paper, we present ReXplain, an innovative system that integrates state-of-the-art modules to produce patient-friendly radiology video reports. Our end-to-end pipeline combines advanced AI capabilities with a practical interface design to address a critical healthcare communication challenge: helping patients better understand their diagnostic results. A proof-of-concept study with six board-certified radiologists yielded encouraging feedback on the utility of these AI-generated video reports. Through the development and initial assessment of ReXplain, we lay the groundwork for AI-assisted, patient-centered radiology reporting.
Appendix A Appendix
Prompt for matching report phrases to organs.
@@ Appendix A Appendix
Prompt for report phrase explanation.
@@
Output: {"abnormality_explanation": ‘‘Your left subclavian vein is found to be collapsed. This vein is an important blood vessel that carries blood from your arm back to your heart. When we say its collapsed, it means that the vein is not as open as it should be, which might affect how blood flows through it.",
"input_scan_appearance": "If we look at your CT image, the left subclavian vein will appear much narrower or even almost closed in some areas. It might look like a thin line instead of a round, hollow tube.",
"normal_scan_appearance": "In a normal CT image, this vein would appear open and round, kind of like a small, hollow tube. It would look uniform and not pinched or narrow."}
List of Organs.
@@
right lung lower lobe’, right lung middle lobe’, right lung upper lobe’,
lung lower lobe’,’lung upper lobe’, trachea and bronchie’, trachea’, bronchie’,
mediastinum’,’brachiocephalic trunk’, brachiocephalic vein’, left brachiocephalic vein’,
@@
spinal cord’, spinal canal’, vertebrae’, cervical vertebrae’,
cervical vertebrae 1 (c1)’, cervical vertebrae 2 (c2)’, cervical vertebrae 3 (c3)’,
cervical vertebrae 4 (c4)’, cervical vertebrae 5 (c5)’, cervical vertebrae 6 (c6)’,
cervical vertebrae 7 (c7)’, thoracic vertebrae’, thoracic vertebrae 1 (t1)’,
thoracic vertebrae 2 (t2)’, thoracic vertebrae 3 (t3)’, thoracic vertebrae 4 (t4)’,
thoracic vertebrae 5 (t5)’, thoracic vertebrae 6 (t6)’,’thoracic vertebrae 7 (t7)’,
thoracic vertebrae 8 (t8)’, thoracic vertebrae 9 (t9)’,’thoracic vertebrae 10 (t10)’,
thoracic vertebrae 11 (t11)’,’thoracic vertebrae 12 (t12)’,’lumbar vertebrae’,
lumbar vertebrae 1 (l1)’, lumbar vertebrae 2 (l2)’,’lumbar vertebrae 3 (l3)’,
lumbar vertebrae 4 (l4)’, lumbar vertebrae 5 (l5)’, lumbar vertebrae 6 (l6)’,
sacral vertebrae 1 (s1)’, clavicle’, left clavicle’, right clavicle’, scapula’,
left scapula’, right scapula’, humerus’, left humerus’, right humerus’, femur’,
left femur’, right femur’, head of femur’, left head of femur’, right head of femur’,
rib’, left rib’, right rib’, rib 1’, left rib 1’, right rib 1’, rib 2’, left rib 2’,
right rib 2’, rib 3’, left rib 3’, right rib 3’, rib 4’, left rib 4’, right rib 4’,
rib 5’, left rib 5’, right rib 5’, rib 6’, left rib 6’, right rib 6’, rib 7’,
left rib 7’, right rib 7’, rib 8’, left rib 8’, right rib 8’, rib 9’, left rib 9’,
right rib 9’, rib 10’, left rib 10’, right rib 10’, rib 11’, left rib 11’,
right rib 11’, rib 12’, left rib 12’, right rib 12’, rib cartilage’, costal cartilage’,
sternum’, manubrium of sternum’, eustachian tube bone’, left eustachian tube bone’,
right eustachian tube bone’, thyroid’, thyroid gland’, left thyroid’, right thyroid’,
breast’, left breast’, right breast’, abdomen’, abdominal tissue’, adrenal gland’,
left adrenal gland’, right adrenal gland’, colon’, duodenum’, gallbladder’, intestine’,
small bowel’, kidney’, left kidney’, right kidney’, liver’, left lobe of liver’,
left lateral inferior segment of liver’, left lateral superior segment of liver’,
left medial segment of liver’, right lobe of liver’,